# Relationship Discovery for Drug Recommendation

### v1

@@
Abstract
Medication recommendation systems are designed to deliver personalized drug suggestions that are closely aligned with individual patient needs. Previous studies have primarily concentrated on developing medication embeddings, achieving significant progress. Nonetheless, these approaches often fall short in accurately reflecting individual patient profiles, mainly due to challenges in distinguishing between various patient conditions and the inability to establish precise correlations between specific conditions and appropriate medications. In response to these issues, we introduce DisMed, a model that focuses on patient conditions to enhance personalization. DisMed employs causal inference to discern clear, quantifiable causal links. It then examines patient conditions in depth, recognizing and adapting to the evolving nuances of these conditions, and mapping them directly to corresponding medications. Additionally, DisMed leverages data from multiple patient visits to propose combinations of medications. Comprehensive testing on real-world datasets demonstrates that DisMed not only improves the customization of patient profiles but also surpasses leading models in both precision and safety.
1 Introduction
@@
Given the noticeable imbalance in healthcare resources’ supply and demand environment, there has been a growing emphasis on AI-based medical systems Rosaci (2007); Choi et al. (2016a); Bansal et al. (2014); Ma et al. (2023). Within this domain, medication recommendation aims to derive treatment strategies from comprehensive clinical samples, utilizing data analysis Ma et al. (2018); Symeonidis et al. (2021); Wang et al. (2021) and deep learning Shang et al. (2018); Wang et al. (2019, 2022d). It then recommends suitable medications for the specific conditions of patients.
Figure 1: Two patients with similar conditions, involving diseases and procedures, receive varying medications by doctor, where red for disease, purple for procedure, and green for medication.
In our study, we observe an intriguing phenomenon, as illustrated in Figure 1, where even though two patients may suffer from similar diseases and procedures, the medications required can be vastly different. Moreover, we conduct a pairwise comparison of medical visit records, and our statistical analysis reveals that among patients with conditions [^1] Patient conditions encompass the health status of individuals, encapsulating both diseases and procedures in healthcare encounter. showing a high similarity of 80%-90%, the average similarity in their prescribed medications was only 48.8%. This observation underscores the importance of personalized patient modeling in the task of medication recommendation.
@@
Previous research methods, such as Yang et al. (2021a); Wu et al. (2022), emphasize the importance of prior medications, suggesting that current medication combinations be recommended based on medications prescribed during the last visit. Meanwhile, other methods, like Yang et al. (2021b, 2023), advocate starting from a molecular perspective to construct more accurate representations of medications. However, these studies are primarily medication-centric and limited to addressing the aforementioned issue of personalization due to the following limitations:
1) Overlooking the difference between patient conditions: The same disease may assume different roles in various clinical visit contexts: it may be the primary cause of other diseases during one visit and a secondary disease resulting from other diseases during another visit. This intricate pathological relationship among diseases implies that the impact of diseases on patients varies across different environments. A similar issue exists with procedures. When dealing with multiple diseases and procedures, doctors need to consider these dynamic factors to provide personalized medication recommendations. However, existing methods uniformly employ equal-weighted summation to represent multiple conditions for all patients, neglecting the differences among conditions.
@@
2) Excessively rough relationship between conditions and medications: Most medications are specifically designed for one or two particular conditions, which means that the relationships between conditions and medications should be point-to-point. Traditional methods fail to discern these specific relationships, typically matching condition sets with medication sets, and constructing set-to-set relationships based on co-occurrence. However, these co-occurrence-based relationships are excessively rough and prone to misleading correlations, which may be progressively amplified in the feedback loop, leading to inaccurate recommendations.
@@
To address the aforementioned limitations, we propose a condition-centric model named DisMed. Specifically, firstly, we incorporate causal relationships to replace the co-occurrence relationships. Through causal discovery, we explore the directional relationships between conditions, and refine the relationships between conditions and medications using causal estimation methods. To address the limitation concerning the differences between conditions, we introduce a novel approach named Dynamic Self-Adaptive Attention (DSA). Grounded in the pathological relationships between conditions, DSA investigates the diverse impacts of conditions on patients across different environments, leading to the construction of personalized patient representations. In response to the limitation of overly rough relationships between conditions and medications, we construct corresponding graph networks by analyzing the quantified therapeutic effects of specific medications on specific diseases. This process enables us to achieve an accurate point-to-point mapping from patient conditions to medications. Ultimately, by applying this method to historical medical visit data, we integrate longitudinal visit information, resulting in tailored medication combination recommendations. We list our main contributions as follows:
  • We present a novel medication recommendation framework focused on the patient’s condition and innovatively introduce causal inference approach in this field to improve personalized patient representation.
  • We observe varying effects on patients based on their different medical conditions and develop a Dynamic Self-Adaptive Attention (DSA) mechanism to capture the dynamic differences in patient conditions.
  • We devise a method that refines the relationship between conditions and medications, transitioning the relationship from set-to-set level to point-to-point level.
@@
  • Through extensive experiments on real-world datasets, we demonstrate our model’s ability to capture personalized patient representations, outperforming other state-of-the-art baselines significantly.
2 Related Work
2.1 Medication Recommendation
Early works in medication recommendation like Zhang et al. (2017) were instance-based, prescribing on single-visit data, and viewing medication recommendation as a multi-instance multi-label classification task. After that, studies like Choi et al. (2016b) employ sequential models and integrating longitudinal visit data. Recently, the research on medication recommendation has significantly increased, with these studies falling into three categories. The first category of work, exemplified by studies like Le et al. (2018); Zheng et al. (2021b); Li et al. (2023), treats medication recommendation as a package, sequence, and other mature recommendation tasks, utilizing advanced algorithms to enhance model quality. The second category focuses on the distinctiveness of medication recommendations. Some studies Yang et al. (2021a); Wu et al. (2022); Zheng et al. (2023) embed medication recommendations within frameworks of translation models, generative models, and other domains, transforming conditions into medication combinations. The third category of work, such as Shang et al. (2019); Yang et al. (2021b, 2023); Chen et al. (2023), incorporates external knowledge related to medications to fill gaps in data, thereby achieving more accurate representations of medications.
@@ 2.1 Medication Recommendation
Our proposed model employs a condition-centric framework and enhances personalized patient representations through the incorporation of causal inference methods.
2.2 Causal Inference in Recommender Systems
@@
The incorporation of causal inference Gao et al. (2022) into recommender systems(RS) still presents as a relatively novel concept, and ongoing research directions can be broadly categorized into three main categories. The primary objective of the first research category is to tackle data bias issues. For instance, methods such as the backdoor criterion employed by Wei et al. (2021); Zheng et al. (2021a); Wang et al. (2022c) aim to eliminate confounding factors, addressing challenges like popularity bias and exposure bias in the dataset. In the second category of work, exemplified by Zhang et al. (2021); Xiao and Wang (2022); Wang et al. (2022b), real-world data collection is coupled with inverse probability weighting and counterfactual methods. This involves introducing counterfactual interactions that contradict observations to fill data gaps or reduce noise. Additionally, a relatively less-explored direction focuses on enhancing model interpretability and fairness. Wang et al. (2022a); Si et al. (2022); Wang et al. (2023) introduce causal discovery and counterfactual approaches in RS to tackle the inherent opacity issues associated with deep learning models.
@@
3 Problem Definition
3.1 Medical Entity
@@
3.2 Input and Output
@@
3.3 DDI Matrix
Our Drug-Drug Interaction (DDI) data is extracted from the Adverse Event Reporting Systems Hoffman et al. (2014). We represent DDI information using a binary matrix ${\mathbf{M}^{ddi}}\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{M}|}$, where $\mathbf{M}^{ddi}_{ij}=1$ indicates the presence of an interaction between $m_{i}$ and $m_{j}$. A high frequency of DDI suggests potential safety issues in recommended results.
3.4 Causal Discovery and Estimation
Causal discovery and estimation constitute two pivotal aspects of causal inference Hasan et al. (2023). Causal discovery entails employing statistical tools and machine learning algorithms to unveil causal relationships between variables. Causal estimation underscores the quantification of causal effects between variables, relying on observed data.
4 The Proposed Model: DisMed
Figure 2: DisMed framework: The upper section illustrates the process of learning the patient’s condition from a single clinical visit. The lower section represents the integration of information from multiple visits and the prediction of medication combinations.
As depicted in Figure 2, our model starts by extracting entity representations from a single clinical visit. In the relationship mining stage, we employ causal discovery and estimation to uncover pathological relationships among conditions and quantify each medication’s impact on these conditions. During the condition analysis phase, we capture dynamic condition variations through causal relationships and map them to the medication level, updating and aggregating embeddings of all medical entities to create a visit representation. Finally, in the information integration stage, we predict medication combinations based on longitudinal patient visit records.
4.1 Entity Representation
@@
$\footnotesize\mathbf{h}_{d_{i}}=\mathbf{E}_{d}(d_{i}),\quad\mathbf{h}_{p_{j}}=\mathbf{E}_{p}(p_{j}),\quad\mathbf{h}_{m_{k}}=\mathbf{E}_{m}(m_{k}),$ (1)
@@
4.2 Relationship Mining
@@
We leverage causal inference to mine causal relationships and effects among medical entities at this stage. Firstly, we analyze the EHR to obtain the distribution $U$ of medical entities and employ a variation of Greedy Intervention Equivalence Search (GIES) Nandy et al. (2018); Kalainathan et al. (2020) to generate a causal relationship graph $G$. This greedy algorithm optimizes Bayesian equivalence class. Concerning the Bayesian equivalence class, we define the scoring criterion as $S(G,U)$ to evaluate the quality of each causal graph $G$ learned from the distribution $U$.
$\footnotesize S(G,U)=\sum_{i=1}^{n}S(X_{i},P^{G}_{a_{i}}),$ (2)
$G^{\prime}=\text{GIES}(S,G),$ (3)
@@
where $n$ is the number of variables, $X_{i}$ is the variable, and $P^{G}_{a_{i}}$ represents the parents of $X_{i}$ in the graph $G$. The $\text{GIES}(\cdot)$ optimizes and learns a optimized graph $G^{\prime}$ from the initial graph $G$ and the equivalence score $S$. We apply the above process to each medical visit, generating a series of pathological graphs $\mathcal{G}=\{G_{v_{1}},G_{v_{2}},\ldots\}$ that encompass all visits.
@@
Subsequently, adopting the principles of causal estimation, we define diseases and procedures as treatment variables and medications as outcome variables. We use a generalized linear model adjusted based on the backdoor criterion Pearl and Mackenzie (2018) to calculate the quantified effects of medications on diseases and procedures. Ultimately, we construct causal effect matrices representing the relationships between diseases and medications, as well as procedures and medications, denoted as $\mathbf{M}^{dm}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{M}|}$ and $\mathbf{M}^{pm}\in\mathbb{R}^{|\mathcal{P}|\times|\mathcal{M}|}$.
4.3 Condition Analysis
In this stage, we analyze differences in patient conditions using causality and refine condition-medication correlations to create personalized visit representations.
4.3.1 Patient Conditions Difference Capturing
In this stage, we delve into the dynamic difference in conditions through causal-based pathological relationships. We extract relationships from the causal graph $G_{v_{t}}$ that exclusively involves entities from the current visit and construct three isomorphic causal graphs: $G^{d}_{v_{t}}$, $G^{p}_{v_{t}}$, and $G^{m}_{v_{t-1}}$.

### v2

# CausalMed: Causality-Based Personalized Medication Recommendation Centered on Patient health state

@@
Abstract.
Medication recommendation systems are developed to recommend suitable medications tailored to specific patient. Previous researches primarily focus on learning medication representations, which have yielded notable advances. However, these methods are limited to capturing personalized patient representations due to the following primary limitations: (i) unable to capture the differences in the impact of diseases/procedures on patients across various patient health states; (ii) fail to model the direct causal relationships between medications and specific health state of patients, resulting in an inability to determine which specific disease each medication is treating. To address these limitations, we propose CausalMed, a patient health state-centric model capable of enhancing the personalization of patient representations. Specifically, CausalMed first captures the causal relationship between diseases/procedures and medications through causal discovery and evaluates their causal effects. Building upon this, CausalMed focuses on analyzing the health state of patients, capturing the dynamic differences of diseases/procedures in different health states of patients, and transforming diseases/procedures into medications on direct causal relationships. Ultimately, CausalMed integrates information from longitudinal visits to recommend medication combinations. Extensive experiments on real-world datasets show that our method learns more personalized patient representation and outperforms state-of-the-art models in accuracy and safety.
Medication Recommendation, Causal Inference, Electronic Health Record
submissionid: 137ccs: Information systems Information retrievalccs: Computing methodologies Causal reasoning and diagnosticsccs: Applied computing Health informatics
1. Introduction
@@
Given the noticeable imbalance in healthcare resources’ supply and demand environment, there has been a growing emphasis on AI-based medical systems (Rosaci, 2007; Choi et al., 2016a; Bansal et al., 2014; Ma et al., 2023). Medication recommendation, as one of the domains, aims to provide personalized medication treatment plans for specific patients using data mining (Ma et al., 2018; Symeonidis et al., 2021; Wang et al., 2021) and deep learning (Shang et al., 2018; Wang et al., 2019, 2022b) methods.
(a) An example: patient 1 has disease $d_{3}$, and Patient 2 has disease $d_{4}$, their health states differ only in one disease, yet the prescribed medications differ significantly.
(b) The relationship between the similarity of the patient’s health state and the similarity of the medications taken by the patients.
Figure 1. The problem of insufficient personalization of medication recommendation.
[…]
The term ”health state” refers to the sum of all diseases and procedures within a single clinical visit. Figure 1 illustrates two real-life cases, explaining why personalized medication recommendation is crucial. In our study, we observed a common phenomenon, as depicted in Figure 1(a), where even if two patients may have highly similar health states, the required medications could be vastly different. Furthermore, we conducted pairwise comparisons of all medical records in MIMIC-III (Johnson et al., 2016), as shown in Figure 1(b), where the results indicate that among patients with medical record similarity ranging from 80% to 90%, the average similarity of prescribed medications was only 48.8%. In conclusion, due to factors such as individual patient differences, doctors often propose different treatment plans for similar diseases/procedures. Hence, personalized representations tailored to each patient are particularly important.
@@
Previous research methods, such as (Yang et al., 2021a; Wu et al., 2022), emphasize the importance of prior medications, suggesting that current medication combinations be recommended based on medications prescribed during the last visit. Meanwhile, other methods, like (Yang et al., 2021b, 2023), advocate starting from a molecular perspective to construct more accurate representations of medications. However, these studies are primarily medication-centric and limited to addressing the aforementioned issue of personalization due to the following limitations:
@@
(1) Failing to model a direct causal relationship between medications and patient’s specific health state: Most medications are specifically designed for one or two particular diseases/procedures, which means that the relationships between diseases/procedures and medications should be point-to-point causal relationships. Traditional methods typically match disease/procedure sets with medication sets, and construct set-to-set relationships based on co-occurrence. However, these co-occurrence-based methods generate excessive false correlations and fail to capture the direct causal relationship between medications and the specific patient health state, making it impossible to determine which specific disease each medication is treating.
(2) Overlooking the difference of diseases in various health states: The same disease may assume different roles in various patient health states: it may be the primary cause of other diseases during one visit and a secondary disease resulting from other diseases during another visit. This intricate pathological relationship among diseases implies that the impact of diseases on patients varies across different health states. A similar issue exists with procedures. In dealing with multiple diseases/procedures, existing methods simply sum them with equal weight, failing to consider the varying roles that diseases play in different health states.
@@
To address the aforementioned limitations, based on causal inference (Hasan et al., 2023), we propose a patient health state-centric model named CausalMed. Specifically, first, to address limitation (1), we replace co-occurrence relationships with causal relationships. Through causal discovery, we eliminate backdoor paths (Gultchin et al., 2020; Li et al., 2022) between medical entities (diseases, procedures, medications). Subsequently, we construct corresponding graph networks by estimating the quantified therapeutic effects of specific medications on specific diseases. Ultimately, we transform the initially chaotic set-to-set relationships into clear point-to-point relationships, modeling the direct causal relationships between medications and specific patient states. At the same time, to address limitation (2), we introduce a novel approach named Dynamic Self-Adaptive Attention (DSA). Grounded in the causal-based pathological relationships of each health state, DSA learns the variability in the effects that medical entities have on patients under different health states, leading to the construction of personalized patient representations. Ultimately, by applying this method to historical medical visit data, we integrate longitudinal visit information, resulting in tailored medication recommendations.
Detailed case study can be found in subsection 5.6. Our source code is publicly available on GitHub [^1] https://github.com/lixiang-222/CausalMed. We list our main contributions as follows:
  • Important Discovery: We make the first discovery that different diseases express significantly different meanings under various patient health states, and current works fail to capture which specific disease each medication is treating. These factors contribute to the low level of personalization in existing medication recommendations.
  • Novel Framework: We propose a novel medication recommendation framework centered on patients’ clinical health state. Utilizing causal inference, we capture point-to-point relationships between diseases and medications and develop a Dynamic Self-Adaptive Attention (DSA) mechanism to discern the dynamic differences in diseases across various health states, thereby enhancing the personalization of patient representations.
@@
  • Extensive Evaluation: Through extensive experiments on real-world datasets, we demonstrate our model’s ability to capture personalized patient representations, outperforming other state-of-the-art baselines significantly.
2. Related Work
2.1. Medication Recommendation
Early works in medication recommendation like LEAP (Zhang et al., 2017) are instance-based, prescribing on single-visit data, and viewing medication recommendation as a multi-instance and multi-label classification task.
The first category of research focuses on modeling the sequential relationships in patients’ longitudinal medical records to address issues of personalization. RETAIN (Choi et al., 2016b) models patients’ histories to personalize disease predictions. MICRON (Yang et al., 2021a) adjusts medications for new symptoms to optimize therapy and minimize side effects. COGNet (Wu et al., 2022) translates patient’s diseases/procedures to medications and reuses effective past treatments.
The second category of methods aims to capture the relationships between entities, thereby addressing the problem of low accuracy in prescriptions. DMNC (Le et al., 2018) combines dynamic memory and deep learning to analyze case data and recommend precise medications. Several studies (Zheng et al., 2021b, 2023; Wu et al., 2023) utilize graph networks to describe the connections between medical entities, enhancing the ability to capture relationships, through innovations in constructing graphs. StratMed(Li et al., 2023) stratifies different disease/procedure-medication relationships, enhancing the representation of sparse relationships to alleviate the long-tail problem in the data.
The third category of methods involves incorporating partial medication-related knowledge to address safety and reliability issues in medication combinations. SafeDrug (Yang et al., 2021b) uses molecular data to improve medication relationship characterization and reduce adverse DDIs. MoleRec (Yang et al., 2023) simulates molecular substructures for better DDI representation. Carmen (Chen et al., 2023) integrates patient history with molecular learning to differentiate similar molecules by function.
@@ 2.1. Medication Recommendation
Our proposed model employs a patient health state-centric framework and enhances personalized patient representations through the incorporation of causal inference methods.
Figure 2. CausalMed framework: The upper section illustrates the process of learning the patient’s representation from a single clinical visit. The lower section represents the integration of information from multiple visits and the prediction of medication combinations.
[…]
2.2. Causal Inference in Recommender Systems
@@
The incorporation of causal inference (Gao et al., 2024) into recommender systems(RS) still presents as a relatively novel concept, and ongoing research directions can be broadly categorized into three main categories. The primary objective of the first research category is to tackle data bias issues. For instance, methods such as the backdoor criterion employed by (Wei et al., 2021; Zheng et al., 2021a; Wang et al., 2022d) aim to eliminate confounding factors, addressing challenges like popularity bias and exposure bias in the dataset.
In the second category of work, exemplified by (Zhang et al., 2021; Xiao and Wang, 2022; Wang et al., 2022c), real-world data collection is coupled with inverse probability weighting and counterfactual methods. This involves introducing counterfactual interactions that contradict observations to fill data gaps or reduce noise.
Additionally, the third category of works focus on enhancing model interpretability and fairness. (Wang et al., 2022a; Si et al., 2022) introduce causal discovery and counterfactual approaches in RS to tackle the inherent opacity issues associated with deep learning models.
@@
3. Problem Definition
3.1. Medical Entity
@@
3.2. DDI Matrix
Our Drug-Drug Interaction (DDI) data is extracted from the Adverse Event Reporting Systems (Hoffman et al., 2014). We represent DDI information using a binary matrix ${\mathbf{M}^{ddi}}\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{M}|}$, where $\mathbf{M}^{ddi}_{ij}=1$ indicates the presence of an interaction between medication $m_{i}$ and medication $m_{j}$. A high frequency of DDI suggests potential safety issues in recommended results.
3.3. Causal Discovery and Estimation
Causal discovery and estimation constitute two pivotal aspects of causal inference (Hasan et al., 2023). Causal discovery entails employing statistical tools and machine learning algorithms to unveil causal relationships between variables. Causal estimation underscores the quantification of causal effects between variables, relying on observed data.
3.4. Input and Output
@@
4. The Proposed Model: CausalMed
As depicted in Figure 2, our model starts by extracting entity representations from a single clinical visit. In the relationship mining stage, we employ causal discovery and estimation to uncover pathological relationships among medical entities and quantify the point-to-point relationships. During the phase of health state learning, we capture the dynamic changes of entities in different patient states through causal relationships, updating and aggregating embeddings of all medical entities to generate visit representations. Finally, in the information integration stage, we predict medication combinations based on longitudinal patient visit records.
4.1. Entity Representation
@@
(1) $\mathbf{h}_{d_{i}}=\mathbf{E}_{d}(d_{i}),\quad\mathbf{h}_{p_{j}}=\mathbf{E}_{p}(p_{j}),\quad\mathbf{h}_{m_{k}}=\mathbf{E}_{m}(m_{k}),$
@@
4.2. Relationship Mining
Figure 3. An example of causal discovery, discovering the true relationship between $d_{1}$ and $d_{2}$ and eliminating the false relationship between $d_{2}$ and $m_{1}$, ultimately transforming set-to-set relationships into point-to-point relationships.
[…]
At this stage, we introduce methods of causal discovery and causal estimation to uncover the underlying genuine relationships among medical entities and eliminate spurious relationships caused by mere co-occurrence. We transform set-to-set relationships into precise point-to-point relationships and refine the causal effects for each group of relationships.
Firstly, as illustrated in Figure 3, the co-occurrence-based method simply associates all diseases and medications that appear together in a single medical visit, assuming that $d_{1}$ and $d_{2}$ are both related to $m_{1}$. However, causal discovery methods can learn that $d_{2}$ is caused by $d_{1}$ by resolving backdoor paths (Pearl and Mackenzie, 2018; Li et al., 2022), and there is no actual relationship between $m_{1}$ and $d_{2}$. We utilize causal discovery to learn the causal structure from set-to-set relationships, uncovering the true underlying relationships and eliminating spurious ones caused by co-occurrence, ultimately generating point-to-point relationships between diseases and medications.
@@ 4.2. Relationship Mining
Specifically, we analyze the EHR to obtain the distribution $U$ of medical entities and employ a variation of Greedy Intervention Equivalence Search (GIES) (Nandy et al., 2018; Kalainathan et al., 2020) to identify confounding bias and generate a causal relationship graph $G$. GIES optimizes the Bayesian equivalence class. Concerning the Bayesian equivalence class, we define the scoring criterion as $S(G,U)$ to evaluate the quality of each causal graph $G$ learned from the distribution $U$.
(2) $S(G,U)=\sum_{i=1}^{n}s(X_{i},P^{G}_{a_{i}}),$
(3) $G^{\prime}=\text{GIES}(S,G),$
@@
where $n$ is the number of variables, $X_{i}$ is the variable including disease, surgery, and medication, and $P^{G}_{a_{i}}$ represents the parents of $X_{i}$ in the graph $G$. And $s(\cdot,\cdot)$ refers to the method of obtaining the equivalence class from $X_{i}$ and its parent nodes ($P^{G}_{a_{i}}$), through which we can produce the initial graph $G$ and continuously update it. The $\text{GIES}(\cdot)$ optimizes and learns an optimized graph $G^{\prime}$ from the initial graph $G$ and the equivalence score $S$. We apply the above process to each medical visit, generating a series of pathological graphs $\mathcal{G}=\{G_{v_{1}},G_{v_{2}},\ldots\}$ that encompass all visits.
@@
Subsequently, adopting the principles of causal estimation, we define diseases and procedures as treatment variables and medications as outcome variables. We use a generalized linear model adjusted based on the backdoor criterion (Pearl and Mackenzie, 2018; Gultchin et al., 2020) to calculate the quantified effects of medications on diseases and procedures. Ultimately, we obtain a quantified relationship between each disease/procedure and each medication, and construct matrices of causal effects, denoted as $\mathbf{M}^{dm}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{M}|}$ and $\mathbf{M}^{pm}\in\mathbb{R}^{|\mathcal{P}|\times|\mathcal{M}|}$.
4.3. Health State Learning
At this stage, we utilize causal relationships to learn the health state of patients. In homomorphic relationships, we capture the meanings of entities under different health states to update their representations; In heterogeneous relationships, we apply point-to-point disease/procedure-medication causality to update entity representations. Ultimately, we integrate this information to create a personalized representation of the clinical visit.
4.3.1. Homomorphic Relationship Learning
As shown in the left half of the Health State Learning part of Figure 2. In this stage, we primarily learn the representations of various entities through the relationships between homomorphic entities. By exploring pathological relationships based on causality, we delve into the dynamic differences between entities to enhance our ability to capture the personalized characteristics of patients.
In different health states, the impact of the same disease on patients can vary. In one state, a disease may be the cause of other diseases, while in another, it may be the result of them. Therefore, we introduce a Dynamic Self-Adaptive Attention (DSA) mechanism to learn the role of diseases in various health states. By using causal positioning to determine the role of a disease in the current clinical setting and assigning different weights, we enhance the personalized representation of patients.